OpenStudy (anonymous):

I need help . Ill give a medal to helper. Write the equation of the line with x-intercept a and y-intercept b.

OpenStudy (anonymous):

a=

OpenStudy (anonymous):

a=4/3 , b= -2

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@annas

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

the formula goes like this x/a + y/b = 1

OpenStudy (anonymous):

I think I must get the LCD first

OpenStudy (anonymous):

oh yes try this one its really simple just put values in the equation

OpenStudy (anonymous):

in intercept form: 3/4x - 2 y = 1 in slope-intercept form: y = -1/2 - 3/2x

OpenStudy (anonymous):

How did you do that? And how it become? 3/4? isnt it 4/3?

OpenStudy (anonymous):

x/4/3 = 3/4 x

OpenStudy (anonymous):

simple law of fractions

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

I am hungry hahahaha. I have some more questions xD gotta grab my breakfast first :D

OpenStudy (anonymous):

xD haha okie bon appetit :)